Determine the numerical value of the following expression without the use of a calculator:

 

log10 (1000100)

100

+

X100

n=1

sin(n) + 1

(􀀀1)n

!



vuut

1Y000

m=1

1

cos(m)2


1
Expert's answer
2021-10-19T12:28:48-0400

Question: Determine the numerical value of the following expression without the use of a calculator:



"(\frac{\log_{10} (1000^{100})} {100} + \sum_{n = 1}^{100}\frac{\sin(\pi n)+1}{(-1)^n}) . \sqrt{\prod_{m=1}^{1000}\frac{1}{\cos (\pi m)^{2}}}"



Solution:

For the logarithm operation:

"\frac{\log_{10} (1000^{100})} {100}"

"= \frac{100\log_{10} 1000} {100}"


"= \frac{100\log_{10} 10^3} {100}"


"= \frac{(3\times 100)\log_{10} 10} {100}"


since "\log_{10} 10 = 1";

Then,



"\therefore\ \frac{\log_{10} (1000^{100})} {100} = \frac{(3\times 100)\log_{10} 10} {100} = \frac{(3\times 100 \times 1)} {100} = 3"


For the cycle operations:

"f(n) = \sum_{n = 1}^{100}\frac{\sin(\pi n)+1}{(-1)^n}"



"for \\ all \\ {n \\in \t\\mathbb{N}}, \\sin(\\pi n) = 0 \\\\ considering \\ the \\ denominator, \\\\ for \\ even \\ n, \\frac{\\sin(\\pi n) + 1} {(-1)^n} =1 \\\\ for \\ odd \\ n, \\frac{\\sin(\\pi n) + 1} {(-1)^n} = -1"

Therefore, dividing the summation into two parts: i.e. 50 odds and 50 evens, we have:

"f(n) = \\ 50 \times (\frac{\sin(\pi n) + 1} {(-1)^n}) \ for\ denominator\ n = evens \\ + 50 \times (\frac{\sin(\pi n) + 1} {(-1)^n}) \ for\ denominator\ n = odds \\\ = (50 \times 1) + (50\times -1) = 50 - 50 = 0"

"\therefore \ f(n) = \sum_{n = 1}^{100}\frac{\sin(\pi n)+1}{(-1)^n} = 0"

Also, for:

"f(m) = \sqrt{\prod_{m=1}^{1000}\frac{1}{\cos (\pi m)^{2}}}"

"for \ all \ {m \in \mathbb{N}}, \frac{1} {\cos(\pi m)^2} = 1"

Therefore the product of:


"\prod_{m=1}^{1000}\frac{1}{\cos (\pi m)^{2}} = 1 \times 1 \times 1 \times........ = 1 \\
\\ \therefore \ the \ square \ root \ of \ the \ product \ of: \\ \sqrt{\prod_{m=1}^{1000}\frac{1}{\cos (\pi m)^{2}}} = \sqrt{1} = 1"


"\therefore \ f(m) = \sqrt{\prod_{m=1}^{1000}\frac{1}{\cos (\pi m)^{2}}} = 1"



Therefore, the final answer to the expression is given by:



"\therefore (\frac{\log_{10} (1000^{100})} {100} + \sum_{n = 1}^{100}\frac{\sin(\pi n)+1}{(-1)^n}) . \sqrt{\prod_{m=1}^{1000}\frac{1}{\cos (\pi m)^{2}}}"




"= (3 + 0) \times 1""= 3 \times 1""= 3"